OpenStudy (anonymous):

Write the equation of the line with a slope of 3/7 passing through the point (9, -2).

OpenStudy (anonymous):

a. 2x-9y=-3/7 b. y=3/7+9 c. y=3/7+2 d. 3x - 7y = 41

OpenStudy (anonymous):

sorry its 3/7x for b and c

OpenStudy (anonymous):

c was y=3/7x-2

OpenStudy (anonymous):

not y=3/7x+2 sorry

OpenStudy (anonymous):

not y=3/7x+2 sorry

OpenStudy (anonymous):

please help